OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center (0, 8), radius = 8

OpenStudy (marissalovescats):

The equation for a circle is: (x-h)^2+(y-k)^2=r^2 Where (h,k) is the center and r^2 is the radius.

OpenStudy (marissalovescats):

You just plug in the numbers you have for (0,8) and radius of 8

OpenStudy (anonymous):

ik but the answer as to be like this __+(___)_______

OpenStudy (marissalovescats):

(x-0)^2+(y-8)^2=8^2 x^2+(y-8)^2=64. When you have a 0 coordinate, it's just x^2 or y^2.

OpenStudy (anonymous):

o thank u sooo much i have one more can u help?

OpenStudy (marissalovescats):

Yeah sure!

OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( -2, -5), radius = 1

OpenStudy (marissalovescats):

This would be the same kind of thing. Plug it in: (x-(-2)^2+(y-(-5)^2=1^2

OpenStudy (marissalovescats):

What happens when you have this -(-)? What does it become?

OpenStudy (anonymous):

a positive?

OpenStudy (marissalovescats):

Good, so what would the equation be?

OpenStudy (anonymous):

(x+2)^2+(y+5)^2=1

OpenStudy (marissalovescats):

Yup!

OpenStudy (anonymous):

omg thank u so much!!!!!! r u good with diagnols?

OpenStudy (marissalovescats):

Hm it depends on the question lol

OpenStudy (anonymous):

A parallelogram whose vertices have coordinates R(1, -1), S(6, 1), T(8, 5), and U(3, 3) has a shorter diagonal of ___ .

OpenStudy (anonymous):

i have 3 answer to chose from

OpenStudy (anonymous):

5 square root(13) square root(97)

OpenStudy (marissalovescats):

Well first thing I would do is graph the parallelogram

OpenStudy (marissalovescats):

 So I drew it and drew the diagonals. So to find the length of the diagonals we will use the distance formula with points

OpenStudy (marissalovescats):

Use the formula with: (3,3)(6,1) and (1,-1)(8,5)

OpenStudy (marissalovescats):

Do you know what the distance formula is/how to use it?

OpenStudy (anonymous):

yup! ill do it right now dont go nowhere though

OpenStudy (marissalovescats):

I won't haha I'll do it too

OpenStudy (anonymous):

haha kk

OpenStudy (anonymous):

so the first one i got square root 13 and the seond square root 185! i dont know how to reduce squareroots

OpenStudy (marissalovescats):

Well it's asking for the shorter diagonal anyways. And for the first one I got sqrt(5)

OpenStudy (anonymous):

really!!? weird

OpenStudy (marissalovescats):

And for the second I got sqrt(85)

OpenStudy (marissalovescats):

Obviously the smaller one is sqrt(5)

OpenStudy (anonymous):

haha ya thank u can u stick around im stuck on alot more!!!!! and u explain really good?

OpenStudy (marissalovescats):

Oops my 1st one is wrong I'm sorry I wrote down (3,3) and it looked like (5,3) lol

OpenStudy (anonymous):

ooo hahaha so u gunna redo it?

OpenStudy (marissalovescats):

It is sqrt(13) lol

OpenStudy (marissalovescats):

Which would be the answer because sqrt(13) is smaller than sqrt(85)

OpenStudy (anonymous):

hahah thanks knew it heheh!!! ur a big help!!! will u help with more?

OpenStudy (marissalovescats):

Sure I can! Open a new question though, this one is getting lengthy lol

OpenStudy (anonymous):

hahaha ok!

OpenStudy (marissalovescats):

Just tag me in your question :)